# Contextual Gesture: Co-Speech Gesture Video Generation through Context-aware Gesture Representation

### v1

@@ Abstract
Co-speech gesture generation is crucial for creating lifelike avatars and enhancing human-computer interactions by synchronizing gestures with speech. Despite recent advancements, existing methods struggle with accurately identifying the rhythmic or semantic triggers from audio for generating contextualized gesture patterns and achieving pixel-level realism. To address these challenges, we introduce Contextual Gesture, a framework that improves co-speech gesture video generation through three innovative components: (1) a chronological speech-gesture alignment that temporally connects two modalities, (2) a contextualized gesture tokenization that incorporate speech context into motion pattern representation through distillation, and (3) a structure-aware refinement module that employs edge connection to link gesture keypoints to improve video generation. Our extensive experiments demonstrate that Contextual Gesture not only produces realistic and speech-aligned gesture videos but also supports long-sequence generation and video gesture editing applications, shown in Fig. 1 Project Page: https://andypinxinliu.github.io/Contextual-Gesture/.
1 Introduction
@@
In human communication, speech is often accompanied by gestures that enhance understanding and convey emotions (De Ruiter et al., 2012). As these non-verbal cues play a vital role in effective interaction (Burgoon et al., 1990), gesture generation plays a key component of natural human-computer interactions: equipping virtual avatars with realistic gesture capabilities become essential in creating immersive interactive experiences.
The relationships between the semantic and emotion content of speech context, the corresponding gestures, and the visual appearance of the speaker’s performance are complex. As such, many recent works (Yi et al., 2023; Liu et al., 2023, 2022d, 2022a) address a reduced form of this problem by generating a simplified representation of the 3D motion, consisting of joints and body parts, that plausibly accompanies a given speech sample, which can then be rendered using standard rendering pipelines. Such representations capture basic motion patterns, yet they neglect the importance of the speaker’s visual appearance, resulting in a lack of realism that hinders effective communication.
Other works, e.g. ANGIE (Liu et al., 2022c) and S2G-diffusion (He et al., 2024), employ image-warping techniques, constrained by keypoints obtained from optical-flow-based deformations, for co-speech video generation. However, such approaches encounter several critical issues. First, these keypoints only define large-scale transformations, and thus miss subtle movements of specific body parts (e.g. hands, fingers). Second, this broad and unconstrained motions representation neglects the triggers associated with gestures, specifically the contextual information from speech. This approach fails to provide the generator with the intrinsic connections between speech context and gesture motion, making it challenging for the generator to produce motion patterns that effectively convey the speaker’s metaphorical expressions or intentions. Finally, the generated motion patterns are often unstructured and overly reactive to large motion, resulting in noisy and imprecise renderings, especially in the hands and shoulders. Collectively, these challenges significantly limit the overall quality and realism of the generated video content.
@@
To address these challenges, we introduce Contextual Gesture, a framework designed to generate speech-aligned gesture motions and high-fidelity speech video outputs. Our approach begins with gesture motion representations using keypoints from human pose estimators. To uncover the intrinsic temporal connections between gestures and speech, we employ chronological contrastive learning to align these two modalities. This joint representation captures the triggers of gesture patterns influenced by speech. We incorporate speech-contextual features into the tokenization process of gesture motions through knowledge distillation, aiming to infuse the gesture representations with implicit intentions conveyed in the audio. This integration creates a clear linkage between the gestures and the corresponding speech, enabling the generation of gestures that reflect the speaker’s intended meaning. For motion generation, we leverage a masking-based gesture generator that refines the alignment of gesture motions with the speech signal through bidirectional mask pretraining. Finally, for uplifting the latent motion generation into 2D animations, we propose a structure-aware image refinement module that generates heatmaps of edge connections from keypoints, providing image-level supervision to improve the quality of body regions with large motion. Extensive experiments demonstrate that our method outperforms the existing approaches in both quantitative and qualitative metrics and achieves long sequence generation and editing applications.
In summary, our primary contributions are:
1.
@@
Contextual Gesture, a framework that achieves joint contextualized gesture motion and video generation with various application support;
2.
@@
a Context-aware gesture representation obtained through knowledge distillation from the chronological gesture-speech aligned features from chronological contrastive learning; and
3.
@@
a Structural-aware Image refinement module, with edge heatmaps as supervision to improve video fidelity.
2 Related Work
Co-speech Gesture generation
@@ 2 Related Work
Most recent works on co-speech gesture generation employ skeleton- or joint-level pose representations. (Ginosar et al., 2019) use an adversarial framework to predict hand and arm poses from audio, and leverage conditional generation (Chan et al., 2019) based on pix2pixHD (Wang et al., 2018) for videos. Some recent works (Liu et al., 2022d; Deichler et al., 2023; Xu et al., 2023; Ao et al., 2022) learns the hierarchical semantics or leverage contrastive learning to obtain joint audio-gesture embedding from linguistic theory to assist the gesture pose generation. TalkShow (Yi et al., 2023) estimates SMPL (Pavlakos et al., 2019) poses, and models the body and hand motions for talk-shows. CaMN (Liu et al., 2022b), EMAGE (Liu et al., 2023) and Diffsheg (Chen et al., 2024) use large conversational and speech datasets for joint face and body modeling with diverse style control. ANGIE (Liu et al., 2022c) and S2G-Diffusion (He et al., 2024) use image-warping features based on MRAA (Siarohin et al., 2021) or TPS (Zhao & Zhang, 2022) to model body motion and achieve speech driven animation by learning correspondence between audio and image-warping features. However, none of these works produce structure- and speech-aware motion patterns suitable for achieving natural and realistic gesture rendering.
Conditional Video Generation
@@
Conditional Video Generation has undergo significant progress for various modalities, like text (Blattmann et al., 2023), pose (Karras et al., 2023; Wang et al., 2023), and audio (Ruan et al., 2023). AnimateDiff (Guo et al., 2024b) presents an efficient low-rank adaptation (Hu et al., 2022) (LoRA) to adapt image diffusion model for video motion generation. AnimateAnyone (Hu et al., 2023) construct referencenet for fine-grained control based on skeleton. Make-Your-Anchor (Huang et al., 2024) and Champ (Zhu et al., 2024) improve avatar video generation through face and body based on SMPL-X conditions. EMO (Tian et al., 2024) and EchoMimic (Meng et al., 2024) leverages audio as control signal for talking head and upper body generation. However, these methods are slow in inference speed and ignore the gesture patterns or rhythmic or semantic signals from audio.
3 Contextual Gesture
@@
Shown in Fig. 2, our framework targets at generating co-speech photo-realistic human videos with contextualized gesture patterns. To achieve this goal, we first learn contextual-aware gesture motion representation through knowledge distillation based on chronological gesture-speech alignment (Sec. 3.1). We leverage a Masking-based Gesture Generator for gesture motion generation. (Sec. 3.2) To improve the noisy hand and shoulder movement during the transfer of latent motion to pixel space, we propose a structure-aware image refinement through edge heatmaps for guidance. (Sec. 3.3).
3.1 Contextualized Gesture Representation
@@
Generating natural and expressive gestures requires capturing fine-grained contextual details that conventional approaches often overlook. Consider a speaker emphasizing the word ”really” in the sentence ”I really mean it” - while existing methods might generate a generic emphatic gesture, they typically fail to capture the subtle, context-specific body movements that make human gestures truly expressive. This limitation stems from relying solely on motion quantization, which often loses the nuanced relationship between speech and corresponding gestures.
@@
To address this challenge, we propose a novel approach that integrates both audio and semantic information into the motion quantizer’s codebook. This integration requires solving two fundamental problems. First, we need to understand how gestures align with speech not just at a high level, but in terms of precise temporal correspondence - when specific words or phrases trigger particular movements, and how the rhythm of speech influences gesture timing. To capture these temporal dynamics, we develop a chronological gesture-speech alignment framework using specialized contrastive learning. Second, we leverage knowledge distillation to incorporate this learned temporal alignment information into the gesture quantizer, enabling our system to generate gestures that are synchronized with speech both semantically and rhythmically.
Feature Representation.
@@
We utilize 2D poses extracted from images to formulate gestures by facial and body movements. We represent a gesture motion sequence as , where denotes the length of the motion, represents the 2D facial landmarks, and denotes the 2D body landmarks. For speech representation, we extract audio embeddings from WavLM (Chen et al., 2022) and Mel spectrogram features (Rabiner & Schafer, 2010) and beat information using librosa (McFee et al., 2015). For text-semantics, we extract embedding from RoBERTa (Liu et al., 2019). These features are concatenated to form the speech representation.
Chronological Speech-Gesture Alignment.
@@
Traditional approaches to modality alignment (Ao et al., 2022; Liu et al., 2022d; Deichler et al., 2023) rely on global representations through class tokens or max pooling, which overlook the fine-grained temporal dynamics between speech and gestures. We address this limitation by introducing chronological modality alignment.
Vanilla Contrastive Alignment.
@@
To align gesture motion patterns with the content of speech and beats, we first project both speech and gesture modalities into a shared embedding space to enhance the speech content awareness of gesture features. As illustrated in Fig. 3 Middle, we separately train two gesture content encoders, for face motion and for body motion, alongside two speech encoders, and , to map face and body movements and speech signals into this joint embedding space. For simplicity, we represent the general gesture motion sequence as . We then apply mean pooling to aggregate content-relevant information to optimize the following loss (van den Oord et al., 2019):
@@
Chronological Negative Examples.
While vanilla Contrastive Learning builds global semantical alignment, we further propose to address the temporal correspondence between speech and gesture. As shown in Fig. 3 Left, consider a speaker saying, ”After watching that video, you realize how much, more than any other president…”. In this case, the gesture sequence involves ”knocking at the table” when saying ”more than any other,” serving as a visual emphasis for ”how much” to highlight the point. To encourage the model understand both semantic and rhythmic alignment between two modalities, we shuffle the words and their corresponding phonemes. By shuffling the sequence to ”you realize how much, after watching that video,” the semantic intention of the speech is preserved, but the rhythmic correspondence between speech and gesture is disrupted. We use Whisper-X (Bain et al., 2023) to detect temporal segments in the raw sequences. We cut the audio and shuffle these segments, creating these augmented samples as additional chronological negative examples within a batch during contrastive learning.
Gesture Quantization with Distillation.
To construct context-aware motion representations, we encode alignment information into the gesture motion codebook. This allows the semantics and contextual triggers from speech to be directly fused into the motion embedding, and enables the generator to easily identify the corresponding motion representation in response to speech triggers. To achieve this goal, we leverage gesture content encoder as the teacher and distill knowledge to codebook latent representation, shown in Fig. 3 Middle. We maximize the cosine similarity over time between the RVQ quantization output and the representation from the gesture content encoder:
@@
3.2 Speech-conditioned Gesture Motion Generation
For Motion Generator, we adopt a similar masking-based generation procedure as in MoMask (Guo et al., 2024a) due to its flexibility of editing and fast inference speed. We only include the generator data flows in the main paper but defer the training and inference strategy in the Appendix.
Mask Gesture Generator.
As shown in Fig. 3 Right, during training, we derive motion tokens by processing raw gesture sequences through both body and face tokenizers. The motion token corresponding to the source image acts as the conditioning for all subsequent frames. For speech control, we initialize the audio content encoder from alignment pre-training as described in Sec. 3.1. This pre-alignment of gesture tokens with audio encoder features enhances the coherence of gesture generation. We employ cross-attention to integrate audio information and apply Adaptive Instance Normalization (AdaIN) (Huang & Belongie, 2017), enabling gesture styles based on the speaker’s identity.
Residual Gesture Generator.
The Residual Gesture Generator shares a similar architecture with the Masked Gesture Generator, but it includes separate embedding layers corresponding to each RVQ residual layer. This module iteratively predicts the residual tokens from the base layers, ultimately producing the final quantized output. Please see MoMask (Guo et al., 2024a) for further details.
3.3 Structure-Aware Image Generation
Converting generated gesture motions into realistic videos presents significant challenges, particularly in handling camera movement commonly found in talkshow and speech videos. While recent 2D skeleton-based animation methods (Hu et al., 2023) offer a potential solution, our empirical analysis (detailed in the Appendix) reveals that these approaches struggle with background stability in the presence of camera motion.
To address this limitation, we draw inspiration from optical-flow based image warping methods (Zhao & Zhang, 2022; Siarohin et al., 2021, 2019), which have shown promise in handling deformable motion. We adapt these approaches by replacing their unsupervised keypoints with our generated gesture keypoints from Sec. 3.2 for more precise foreground optical flow estimation. However, this estimation still presents uncertainties, resulting in blurry artifacts around hands and shoulders, particularly when speakers make rapid or large movements.

### v2

@@ Abstract.
Co-speech gesture generation is crucial for creating lifelike avatars and enhancing human-computer interactions by synchronizing gestures with speech. Despite recent advancements, existing methods struggle with accurately identifying the rhythmic or semantic triggers from audio for generating contextualized gesture patterns and achieving pixel-level realism. To address these challenges, we introduce Contextual Gesture, a framework that improves co-speech gesture video generation through three innovative components: (1) a chronological speech-gesture alignment that temporally connects two modalities, (2) a contextualized gesture tokenization that incorporate speech context into motion pattern representation through distillation, and (3) a structure-aware refinement module that employs edge connection to link gesture keypoints to improve video generation. Our extensive experiments demonstrate that Contextual Gesture not only produces realistic and speech-aligned gesture videos but also supports long-sequence generation and video gesture editing applications, shown in Fig. 1.
1. Introduction
@@
In human communication, speech is often accompanied by gestures that enhance understanding and convey emotions (De Ruiter et al., 2012). These non-verbal cues play a crucial role in effective interaction (Burgoon et al., 1990), making gesture generation a key component of natural human-computer interaction. Equipping virtual avatars (Song et al., 2024a; Liu et al., 2024a; Zhang et al., 2025) with realistic gesture capabilities is therefore essential for creating engaging interactive experiences for industry production (Tang et al., 2025).
Since the era of deep learning paradigm (Zhu et al., 2024b; Lu et al., 2022; Gao et al., 2025; Lu et al., 2023, 2024; Ning et al., 2024), many recent works tried to model the relationship between the semantic and emotional content of speech, the associated gestures (Yi et al., 2023; Liu et al., 2023; Luo et al., 2022; Liu et al., 2022c, a, 2025b). They simplify the problem by generating coarse 3D motion representations—typically keypoints of joints and body parts—that plausibly align with a given speech sample. These simplified representations can be rendered via standard pipelines and effectively capture basic motion patterns. However, they often disregard the speaker’s visual appearance and fine-grained motions, leading to a lack of realism that limits the expressive power and communicative effectiveness of the generated gestures.
Other works, such as ANGIE (Liu et al., 2022b) and S2G-diffusion (He et al., 2024), employ image-warping techniques for video generation, guided by motion representations in the form of keypoints obtained from optical-flow-based deformations. While promising, these approaches face several critical limitations. First, their broad and unconstrained motion representations fail to capture the contextual triggers of gestures—particularly the semantic and emotional cues embedded in speech. This disconnect limits the model’s ability to learn the nuanced relationship between speech and gesture, making it difficult to generate expressive motions that convey the speaker’s intent or metaphoric meaning. Second, the predicted keypoints primarily encode coarse, large-scale transformations, resulting in unstructured motion patterns that are overly sensitive to pronounced movements. Consequently, the generated outputs often suffer from noise and imprecision, especially in fine-grained areas such as the hands and shoulders. These issues degrade the realism and coherence of the generated video content.
@@
To address these challenges, we introduce Contextual Gesture, a framework designed to generate speech-aligned gesture motions and afterwards high-fidelity speech video. To uncover the intrinsic temporal connections between gestures and speech, we employ chronological contrastive learning to align these two modalities into time-sensitive joint representation of speech-contextual and gesture motion, which captures the triggers of gesture patterns influenced by speech. We then use knowledge distillation to incorporate speech-contextual features into the tokenization process of gesture motions, aiming to infuse the implicit intentions of gestures conveyed in the speech. This integration creates a clear chronological linkage between the gestures and the speech, enabling the generation of gestures that reflect the speaker's intentions. For motion generation, we employ a masking-based, Transformer gesture generator that produces motion tokens and refines their alignment with the speech signal through bidirectional mask pretraining. Finally, for uplifting the gesture generation into 2D animations, we propose a structure-aware image refinement module that generates heatmaps of edge connections from keypoints, providing image-level supervision to improve the quality of body regions with large motion. Extensive experiments demonstrate that our method outperforms the existing approaches in both quantitative and qualitative metrics and achieves long sequence generation and editing applications. In summary, our primary contributions are:
@@
(1) Contextual Gesture, a framework that achieves time-sensitive joint representation of both speech-contextual and gesture motion, and video generation with various application support;
@@
(2) a Contextual-aware gesture representation obtained through knowledge distillation from the chronological gesture-speech aligned features from chronological contrastive learning;
@@
(3) a Structural-aware Image refinement module, with edge heatmaps as supervision to improve video fidelity.
2. Related Work
@@
Co-speech Gesture generation Co-speech gesture generation is essential for digital avatar production (Song et al., 2024b; Kumar et al., 2023) and functions as the prior for realistic rendering (Luo et al., 2024; Song et al., 2024c, 2023). (Ginosar et al., 2019) use an adversarial framework to predict hand and arm poses from audio, and leverage conditional generation (Chan et al., 2019) based on pix2pixHD (Wang et al., 2018) for videos. Some recent works (Liu et al., 2022c; Deichler et al., 2023; Xu et al., 2023; Ao et al., 2022; Song et al., 2021a, b; Liu et al., 2025a; Song et al., 2024d) learns the hierarchical semantics or leverages contrastive learning to obtain joint audio-gesture embedding from linguistic theory to assist the gesture pose generation. TalkShow (Yi et al., 2023) estimates SMPL (Pavlakos et al., 2019) poses, and models the body and hand motions for talk-shows. CaMN (Liu et al., 2022d), EMAGE (Liu et al., 2023) and Diffsheg (Chen et al., 2024) use large conversational and speech datasets for joint face and body modeling with diverse style control. ANGIE (Liu et al., 2022b) and S2G-Diffusion (He et al., 2024) use image-warping features based on MRAA (Siarohin et al., 2021) or TPS (Zhao and Zhang, 2022) to model body motion and achieve speech driven animation by learning correspondence between audio and image-warping features. However, none of these works produce structure- and speech-aware motion patterns suitable for achieving natural and realistic gesture rendering.
@@
Conditional Video Generation Conditional Video Generation has undergo significant progress with diffusion models (Huang et al., 2025; Tang et al., 2024a; Karras et al., 2023; Huang et al., 2024a; Tang et al., 2024b). AnimateDiff (Guo et al., 2024b) presents an efficient low-rank adaptation (Hu et al., 2022) (LoRA) to adapt image diffusion model for video motion generation. AnimateAnyone (Hu et al., 2023) construct referencenet for fine-grained control based on skeleton. Make-Your-Anchor (Huang et al., 2024b) and Champ (Zhu et al., 2024a) improve avatar video generation through face and body based on SMPL-X conditions. EMO (Tian et al., 2024) and EchoMimic (Meng et al., 2024) and Tango (Liu et al., 2024b) leverages audio as control signal for talking head and upper body generation. However, these methods are slow in inference speed and ignore the gesture patterns or rhythmic or semantic signals from audio.
3. Contextual Gesture
@@
Shown in Fig. 2, our framework targets at generating co-speech photo-realistic human videos with contextualized gestures. To achieve this goal, we first learn time-sensitive contextual-aware gesture representation through knowledge distillation based on chronological gesture-speech alignment (Sec. 3.1). We then leverage a Masking-based Gesture Generator for gesture motion generation. (Sec. 3.2) To improve the noisy hand and shoulder movement during the transfer of latent motion to pixel space, we propose a structure-aware image refinement through edge heatmaps for guidance. (Sec. 3.3).
3.1. Contextualized Gesture Representation
@@
Generating natural and expressive gestures requires capturing fine-grained contextual details that conventional approaches often overlook. Consider a speaker emphasizing the word "really" in the sentence "I really mean it" - while existing methods might generate a generic emphatic gesture, they typically fail to capture the subtle, context-specific body movements that make human gestures truly expressive. This limitation stems from relying solely on motion quantization, which often loses the nuanced relationship between speech and corresponding gestures.
@@
To address this challenge, we propose a novel approach that integrates both audio and semantic information into the motion quantizer's codebook. This integration requires solving two fundamental problems. First, we need to understand how gestures align with speech not just at a high level, but in terms of precise temporal correspondence - when specific words or phrases trigger particular movements, and how the rhythm of speech influences gesture timing. To capture these temporal dynamics, we develop a chronological gesture-speech alignment framework using specialized contrastive learning. Second, we leverage knowledge distillation to incorporate this learned temporal alignment information into the gesture quantizer, enabling our system to generate gestures that are synchronized with speech both semantically and rhythmically.
@@
Feature Representation. We utilize 2D poses extracted from images to formulate gestures by facial and body movements. We represent a gesture motion sequence as , where denotes the length of the motion, represents the 2D facial landmarks, and denotes the 2D body landmarks. For speech representation, we extract audio embeddings from WavLM (Chen et al., 2022) and Mel spectrogram features (Rabiner and Schafer, 2010) and beat information using librosa (McFee et al., 2015). For text-semantics, we extract embedding from RoBERTa (Liu et al., 2019). These features are concatenated to form the speech representation.
@@
Chronological Speech-Gesture Alignment. Traditional approaches to modality alignment (Ao et al., 2022; Liu et al., 2022c; Deichler et al., 2023) rely on global representations through class tokens or max pooling, which overlook the fine-grained temporal dynamics between speech and gestures. We address this limitation by introducing chronological modality alignment.
@@
Vanilla Contrastive Alignment. To align gesture motion patterns with the content of speech and beats, we first project both speech and gesture modalities into a shared embedding space to enhance the speech content awareness of gesture features. As illustrated in Fig. 3 Middle, we separately train two gesture content encoders, for face motion and for body motion, alongside two speech encoders, and , to map face and body movements and speech signals into this joint embedding space. For simplicity, we represent the general gesture motion sequence as . We then apply mean pooling to aggregate content-relevant information to optimize the following loss (van den Oord et al., 2019):
@@
Chronological Negative Examples. While vanilla Contrastive Learning builds global semantical alignment, we further propose to address the temporal correspondence between speech and gesture. As shown in fig. 3 Left, consider a speaker saying, "After watching that video, you realize how much, more than any other president…". In this case, the gesture sequence involves "knocking at the table" when saying "more than any other," serving as a visual emphasis for "how much" to highlight the point. To encourage the model understand both semantic and rhythmic alignment between two modalities, we shuffle the words and their corresponding phonemes. By shuffling the sequence to "you realize how much, after watching that video," the semantic intention of the speech is preserved, but the rhythmic correspondence between speech and gesture is disrupted. We use Whisper-X (Bain et al., 2023) to detect temporal segments in the raw sequences. We cut the audio and shuffle these segments, creating these augmented samples as additional chronological negative examples within a batch during contrastive learning.
Gesture Quantization with Distillation. To construct context-aware motion representations, we encode alignment information into the gesture motion codebook. This allows the semantics and contextual triggers from speech to be directly fused into the motion embedding, and enables the generator to easily identify the corresponding motion representation in response to speech triggers. To achieve this goal, we leverage gesture content encoder as the teacher and distill knowledge to codebook latent representation, shown in Fig. 3 Middle. We maximize the cosine similarity over time between the RVQ quantization output and the representation from the gesture content encoder:
@@
3.2. Speech-conditioned Gesture Generation
For Motion Generator, we adopt a similar masking-based generation procedure as in MoMask (Guo et al., 2024a). We only include the generator data flows in the main paper but defer the training and inference strategy in the Appendix.
Mask Gesture Generator. As shown in Fig. 3 Right, during training, we derive gesture tokens by processing raw gesture sequences through both body and face tokenizers. The gesture token corresponding to the source image acts as the conditioning for all subsequent frames. For speech control, we initialize the audio content encoder from alignment pre-training as described in Sec. 3. This pre-alignment of gesture tokens with audio encoder features enhances the coherence of gesture generation. We employ cross-attention to integrate audio information and apply Adaptive Instance Normalization (AdaIN) (Huang and Belongie, 2017), enabling gesture styles based on the speaker's identity.
Residual Gesture Generator. The Residual Gesture Generator shares a similar architecture with the Masked Gesture Generator, but it includes separate embedding layers corresponding to each RVQ residual layer. This module iteratively predicts the residual tokens from the base layers, ultimately producing the final quantized output. Please see MoMask (Guo et al., 2024a) for further details.
3.3. Structure-Aware Image Generation
Converting generated gesture motions into realistic videos presents significant challenges, particularly in handling camera movement commonly found in videos. While recent 2D skeleton-based animation methods (Hu et al., 2023) offer a potential solution, our empirical analysis (detailed in the Appendix) reveals that these approaches struggle with background stability in the presence of camera motion.
To address this limitation, we draw inspiration from optical-flow based image warping methods (Zhao and Zhang, 2022; Siarohin et al., 2021, 2019), which have shown promise in handling deformable motion. We adapt these approaches by replacing their unsupervised keypoints with our generated gesture keypoints from Sec. 3.2 for more precise foreground optical flow estimation. However, this estimation still presents uncertainties, resulting in blurry artifacts around hands and shoulders, particularly when speakers make rapid or large movements.